# Enhanced Generative Recommendation via Content and Collaboration Integration

### v1

@@ 5.4. Implementation Details
Table 4. Ablation study on three datasets. The best and the second-best scores are marked in bold and underlined fonts, respectively. R@{5,10} and N@{5,10} stand for Recall@{5,10} and NDCG@{5,10} respectively.
Beauty Sports Phone
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
ColaRec 0.0667 0.0993 0.0449 0.0556 0.0442 0.0660 0.0294 0.0364 0.0745 0.1121 0.0490 0.0611
(1) w/o textual content 0.0527 0.0809 0.0346 0.0439 0.0364 0.0542 0.0239 0.0297 0.0636 0.0974 0.0426 0.0535
(2) w/o indexing 0.0637 0.0947 0.0428 0.0531 0.0422 0.0644 0.0278 0.0350 0.0728 0.1086 0.0487 0.0602
(3) w/o $\mathcal{L}_{bpr}$ 0.0612 0.0918 0.0412 0.0512 0.0424 0.0634 0.0282 0.0350 0.0719 0.1077 0.0486 0.0602
(4) w/o $\mathcal{L}_{c}$ 0.0657 0.0989 0.0434 0.0544 0.0422 0.0649 0.0279 0.0352 0.0731 0.1106 0.0485 0.0606
Table 5. Comparison between different types of GIDs. The best and the second-best scores are marked in bold and underlined fonts, respectively. R@{5,10} and N@{5,10} stand for Recall@{5,10} and NDCG@{5,10} respectively.
Beauty Sports Phone
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
ColaRec 0.0667 0.0993 0.0449 0.0556 0.0442 0.0660 0.0294 0.0364 0.0745 0.1121 0.0490 0.0611
(1) $iad$ 0.0658 0.0983 0.0437 0.0544 0.0428 0.0657 0.0285 0.0359 0.0719 0.1074 0.0474 0.0589
(2) $Random$ 0.0600 0.0902 0.0401 0.0500 0.0411 0.0623 0.0272 0.0340 0.0667 0.1004 0.0443 0.0551
(3) $Content$ 0.0662 0.1008 0.0440 0.0554 0.0423 0.0647 0.0278 0.0350 0.0716 0.1080 0.0477 0.0593
@@ 5.5.1. Comparison on whole users
Table 2 shows the performance comparison on overall users. From these results, we make the following observations. Firstly, the proposed ColaRec achieves the best recommendation performance on all datasets except for NDCG@5 in Beauty, which achieves comparable scores with the best NCL baseline. In particular, ColaRec consistently outperforms previous CF-based and generative baselines in Recall@20, achieving a relative improvement of 3.87%, 6.17% and 6.72% on Beauty, Sports, and Phone, respectively. These results demonstrate the effectiveness of ColaRec and its generalization across different domains.
Secondly, compared with DSI-R and DSI-S, which directly adapt generative retrieval methods to the recommendation task, ColaRec achieves a notable 47.89%, 38.13%, and 44.70% relative improvement in terms of Recall@5 on three datasets respectively, demonstrating effectiveness of the proposed ColaRec. Such results also demonstrate that naively transferring the generative retrieval methods for recommendation cannot achieve satisfying results. Furthermore, our method consistently outperforms the state-of-the-art generative recommendation model TIGER, which uses a RQ-VAE module to construct GIDs. The reason may be that TIGER only considers item content information but the collaborative signals are overlooked without an explicit alignment process.
@@ 5.5.2. Comparison on long-tail users
We also conduct experiments to verify the recommendation performance of ColaRec on long-tail users with sparse interactions. In this experiment, the ratio between head users and long-tail users is set as 20%:80%. Table 3 reports the results of Recall on three datasets. Results of NDCG show similar trends and are omitted due to the reason of space. We can see that ColaRec significantly outperforms all baselines when generating recommendation for long-tail users. The reason is that ColaRec models both user-item interactions and item content information. Given the long-tail users with less interaction knowledge, the content information helps ColaRec to gain better performance.
@@ 5.6. Ablation Study (RQ2)
In this section, we conduct ablation studies to analyse the effectiveness of each component in ColaRec. We implement four ablative variants of ColaRec, including: (1) w/o textual contentdeletes all textual content information in the model input; (2) w/o indexingremoves the item-item indexing task; (3) w/o $\mathcal{L}_{bpr}$removes the ranking loss $\mathcal{L}_{\text{bpr}}$; and (4) w/o $\mathcal{L}_{c}$removes the contrastive loss $\mathcal{L}_{\text{c}}$.
The results on three datasets are shown in Table 4. These results clearly indicate that the removal of any component from our proposed method results in a noticeable decline in overall performance. From the results of variant (1), we see that removing textual content information leads to a significant reduction in performance. Specifically, a $26.57\%$, $21.43\%$, and $17.14\%$ drop in Recall@5 is observed across the three datasets. This suggests the importance of incorporating textual content information to enhance the models’ understanding of items. Furthermore, the variant (3), which is trained without $\mathcal{L}_{\text{bpr}}$, exhibits a notable performance decrease in comparison to ColaRec. This highlights the effectiveness of the pairwise ranking objective, which focuses on the relative prioritization of positive and negative items within the generative recommendation paradigm. Besides, the results of variant (2) and (4) verify the effectiveness of the proposed item-item indexing task and the contrastive objective $\mathcal{L}_{c}$. A consistent performance reduction is observed on the three datasets when either of the two techniques is removed. This suggests that aligning item content information with user-item collaborative signals not only facilitates mutual reinforcement but also enables the learning of more comprehensive and effective representations.
@@
(c) Phone
Figure 4. Impact of the length of GIDs. The GID length $l$ is set from $1$ to $4$.
@@
(c) Phone
Figure 5. Impact of the number of clusters. The number of clusters $K$ varies from $32$ to $128$.
@@ 5.7. GID Investigation (RQ3)
The construction of GIDs plays a crucial role in generative recommendation as it defines the model’s search space for generation. To answer RQ3, we conduct two analytical experiments about GIDs: (1) an analysis of different GID types (e.g., the single token $iad$, the random assigned GID, and the content-based GID), and (2) an analysis of hyper-parameters when constructing GIDs (e.g., the length and $K$-means cluster numbers).
@@ 5.7.1. Effect of different GID types
To further evaluate our GID generation strategy, we conducted an ablation study comparing it with three techniques: iad-based GID, $Random$ GID, and $Content$ GID. Specifically, the iad-based GID assigns a unique single token to each item to represent it using the corresponding vector in the item embedding matrix. The Random GID assigns a random string to each item as the identifier, without considering any prior knowledge. Content GID represents the GID constructed from item content information. Unlike the collaborative GID used in ColaRec, the content-based GID employs a hierarchical $K$-means clustering algorithm (Tay et al., 2022) to cluster items based on their textual representation derived from a pretrained BERT. To make a fair comparison with ColaRec, the length and the codebook size of both $Random$ GIDs and $Content$ GIDs are identical to those in ColaRec.
Table 5 details our results on the three datasets, where ColaRec attains the best performance. This indicates that our GID construction method, based on collaborative signals, is highly effective for generative recommendation. The improvement over the Content GID highlights the effectiveness of collaborative signals in recommendation tasks. Furthermore, ColaRec’s superior performance compared to the iad method demonstrates the benefit of explicitly introducing item correlations into item GIDs. In addition, the Random method leads to the lowest performance as the random string could further introduce noisy signals in the learning process. These results illustrate the importance of constructing effective GIDs for generative recommenders.
@@ 5.7.2. Impact of GID Hyper-parameters
In this section, we investigate the impact of hyperparameters $l$ and $K$ in the GID construction process. Figure 4 shows the results of ColaRec with different GID lengths $l$, ranging from $1$ to $4$. We can see that as $l$ varies, the recommedation performance exhibits some fluctuations, with the best results typically observed when $l=3$. It is worth noting that the GID of ColaRec consists of two parts: (i) the prefix tokens $[z_{i}^{1},\dots,z_{i}^{l-1}]$ from hierarchy clustering; and (ii) the last token $z_{l}$, which is randomly assigned to ensure uniqueness of the GID. A longer GID tends to include a greater proportion of hierarchy clustering parts, potentially facilitating model learning. However, a too long GID means more autoregressive decoding steps in model generation, which increases search difficulty and inference latency. Therefore, balancing efficiency and effectiveness, we have selected $l=3$ as the default setting.
To assess how the number of clusters $K$ affects performance, we fixed the GID length $l$ as $l=3$ and varied $K$ with values $32$, $64$, $96$, and 128. As shown in Figure 5, a higher $K$ typically results in a slight decrease in overall performance. The decrease on Beauty is more notable. The reason is that a higher $K$ indicates a larger search space in the decoding process, and thus increase the generation difficulty. To this end, choosing a suitable clustering number according to number of items is important for generative recommendation.
@@ 6. Conclusions and future work
This paper proposes ColaRec, a novel framework to conduct content-based collaborative generation for recommender systems. As an end-to-end generative recommender, ColaRec effectively integrates both item content information and user-item collaborative signals within a unified framework through fine-tuning an encoder-decoder based language model. In addition, an auxiliary item indexing task and a contrastive loss are proposed to further align the model’s representations of item content and user-item collaborative signals. We have conducted extensive experiments on three real-world datasets and empirical results demonstrate the effectiveness of ColaRec.
In the future, we intend to investigate more methods to construct appropriate generative identifiers (GIDs) and adopt more effective approaches to better align content information and collaborative signals. We also plan to introduce larger language models together with larger volume of training data to generate better recommendation. Negative sampling for generative recommendation is one of our future works, either using GIDs to sample more informative negative samples or utilizing generative models to generate simulated negative instances. Besides, parameter efficient fine-tuning and model efficiency are also potential research directions.

### v2

@@ 5.5.1. Comparison on whole users
Table 2 shows the performance comparison of overall users. From these results, we make the following observations. Firstly, ColaRec achieves the best recommendation performance on all datasets except for NDCG@5 in Beauty, which achieves comparable scores with the best NCL baseline. In particular, ColaRec consistently outperforms previous CF-based and generative models in Recall@20, achieving a relative improvement of 3.87%, 3.54%, 6.72% and 3.88% on Beauty, Sports, Phone and Recipe, respectively. These results demonstrate the effectiveness of ColaRec and its generalization across different domains.
Secondly, compared with DSI-R and DSI-S, which directly adapt generative retrieval methods to the recommendation task, ColaRec achieves a notable 47.89%, 38.13%, 80.82% and 26.11% relative improvement in terms of Recall@5 on four datasets respectively, demonstrating the effectiveness of the proposed ColaRec. Such results also demonstrate that naively transferring the generative retrieval methods for recommendation cannot achieve satisfying results. Furthermore, our method consistently outperforms existing representative generative methods TIGER and LC-Rec. For TIGER, the reason may be that it only considers item content information but the collaborative signals are overlooked without an explicit alignment process. While for LC-Rec, the recommendation is conducted through language modeling and generation under designed prompts, which is not fully aligned with the recommendation task, leading to sub-optimal results.
@@
Table 3. Performance comparison of long-tail users. The best and the second-best scores are marked in bold and underlined fonts, respectively. ** denotes the improvements are significant with p-value ¡ 0.05. R stands for Recall.
CF-based Methods Generative Models for Recommendation
Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER LC-Rec Ours
Beauty R@5 0.0416 0.0636 0.0555 0.0639 0.0510 0.0464 0.0131 0.0415 0.0487 0.0492 0.0660**
R@10 0.0604 0.0922 0.0825 0.0907 0.0742 0.0662 0.0228 0.0653 0.0745 0.0772 0.0975**
R@20 0.0817 0.1253 0.1160 0.1264 0.1039 0.0917 0.0354 0.0940 0.1084 0.1107 0.1327**
Sports R@5 0.0209 0.0433 0.0355 0.0440 0.0329 0.0267 0.0116 0.0307 0.0380 0.0397 0.0456**
R@10 0.0317 0.0639 0.0562 0.0645 0.0495 0.0394 0.0170 0.0472 0.0581 0.0617 0.0674**
R@20 0.0468 0.0904 0.0836 0.0908 0.0725 0.0553 0.0273 0.0728 0.0882 0.0929 0.0976**
Phone R@5 0.0405 0.0723 0.0660 0.0727 0.0571 0.0451 0.0206 0.0404 0.0602 0.0612 0.0756**
R@10 0.0590 0.1054 0.0986 0.1043 0.0861 0.0641 0.0371 0.0623 0.0898 0.0918 0.1131**
R@20 0.0855 0.1482 0.1418 0.1473 0.1228 0.0899 0.0600 0.0939 0.1293 0.1353 0.1590**
Recipe R@5 0.0128 0.0204 0.0121 0.0210 0.0182 0.0172 0.0157 0.0171 0.0181 0.0189 0.0219**
R@10 0.0229 0.0320 0.0212 0.0322 0.0309 0.0269 0.0274 0.0295 0.0316 0.0313 0.0334**
R@20 0.0371 0.0487 0.0343 0.0490 0.0499 0.0412 0.0443 0.0475 0.0504 0.0493 0.0528**
@@ 5.5.2. Comparison on long-tail users
We also conduct experiments to verify the recommendation performance of ColaRec on long-tail users with sparse interactions. In this experiment, the ratio between head users and long-tail users is set as 20%:80%. Table 3 reports the results of Recall on four datasets. Results of NDCG show similar trends and are omitted due to the reason of space. We can see that ColaRec significantly outperforms all baselines when generating recommendation for long-tail users. The reason is that ColaRec models both user-item interactions and item content information. Given the long-tail users with less interaction knowledge, ColaRec gains better performance with the help of the content information.
@@
Table 4. Effect of multi-task learning and different types of GIDs. The best scores are marked in bold.
Beauty Sports Phone Recipe
Recall@5 NDCG@5 Recall@5 NDCG@5 Recall@5 NDCG@5 Recall@5 NDCG@5
ColaRec 0.0667 0.0449 0.0442 0.0294 0.0745 0.0490 0.0198 0.0151
(1) w/o textual content 0.0527 0.0346 0.0364 0.0239 0.0636 0.0426 0.0181 0.0141
(2) w/o indexing 0.0637 0.0428 0.0422 0.0278 0.0728 0.0487 0.0179 0.0142
(3) w/o $\mathcal{L}_{bpr}$ 0.0612 0.0412 0.0424 0.0282 0.0719 0.0486 0.0184 0.0140
(4) w/o $\mathcal{L}_{c}$ 0.0657 0.0434 0.0422 0.0279 0.0731 0.0485 0.0188 0.0145
(1) $iad$ 0.0658 0.0437 0.0428 0.0285 0.0719 0.0474 0.0189 0.0145
(2) $Random$ 0.0600 0.0401 0.0411 0.0272 0.0667 0.0443 0.0190 0.0149
(3) $Content$ 0.0662 0.0440 0.0423 0.0278 0.0716 0.0477 0.0183 0.0141
@@ 5.6. Ablation Study (RQ2)
In this section, we conduct ablation studies to analyse the effectiveness of each component in ColaRec. We implement four ablative variants of ColaRec, including: (1) w/o textual contentdeletes all textual content information in the model input; (2) w/o indexingremoves the item-item indexing task; (3) w/o $\mathcal{L}_{\text{bpr}}$removes the ranking loss $\mathcal{L}_{\text{bpr}}$; and (4) w/o $\mathcal{L}_{c}$removes the contrastive loss $\mathcal{L}_{\text{c}}$.
The results are shown in the upper part of Table 4. These results clearly indicate that the removal of any component from our proposed method results in a noticeable decline in overall performance. From the results of variant (1), we see that removing textual content information leads to the significant performance downgrade. Specifically, a $20.99\%$, $17.65\%$, $14.63\%$ and $8.59\%$ drop in Recall@5 is observed across the four datasets, respectively. This suggests the importance of incorporating textual content information to enhance the models’ understanding of items. Furthermore, the variant (3), which is trained without $\mathcal{L}_{\text{bpr}}$, exhibits a notable performance decrease compared to ColaRec. This highlights the effectiveness of the pairwise ranking objective, which focuses on the relative prioritization of positive items within the generative recommendation paradigm. Besides, the results of variant (2) and (4) verify the effectiveness of the proposed item-item indexing task and the contrastive objective $\mathcal{L}_{c}$. A consistent performance reduction is observed on the four datasets when either of the two techniques is removed. This suggests that aligning item content information with user-item collaborative signals not only facilitates mutual reinforcement but also enables the learning of more comprehensive and effective representations.
@@
Figure 4. Impact of the length of GIDs.
@@
Figure 5. Impact of the number of clusters.
@@ 5.7. GID Investigation (RQ3)
The construction of GIDs plays a crucial role in generative recommendation as it defines the model’s search space for generation. To answer RQ3, we conduct two analytical experiments about GIDs: (1) an analysis of different GID types (e.g., the single token $iad$, the random assigned GID, and the content-based GID), and (2) an analysis of hyper-parameters when constructing GIDs (e.g., the length and the $K$-means cluster number).
@@ 5.7.1. Effect of different GID types
To further evaluate our GID generation strategy, we conduct an ablation study comparing it with three techniques: iad-based GID, $Random$ GID, and $Content$ GID. Specifically, the iad-based GID assigns a unique single token to each item to represent it using the corresponding vector in the item embedding matrix. The Random GID assigns a random string to each item as the identifier, without considering any prior knowledge. The Content GID represents the GID constructed from item content information. Unlike the collaborative GID used in ColaRec, the content-based GID employs a hierarchical $K$-means clustering algorithm (Tay et al., 2022) to cluster items based on their textual representation derived from a pretrained BERT model. To make a fair comparison with ColaRec, the length and the codebook size of both $Random$ GIDs and $Content$ GIDs are identical to those in ColaRec.
The bottom part of Table 4 details our results on the four datasets, where ColaRec attains the best performance. This indicates that our GID construction method, based on collaborative signals, is highly effective for generative recommendation. The improvement over the Content GID highlights the effectiveness of collaborative signals in recommendation tasks. Furthermore, ColaRec’s superior performance compared to the iad method demonstrates the benefit of explicitly introducing item correlations into item GIDs. In addition, the Random method leads to the lowest performance as the random string could further introduce noisy signals in the learning process. These results illustrate the importance of constructing effective GIDs for generative recommenders.
@@ 5.7.2. Impact of GID Hyper-parameters
In this section, we investigate the impact of hyper-parameters $l$ and $K$ in the GID construction process. Figure 4 shows the results of ColaRec with different GID lengths $l$, ranging from $1$ to $4$. Note that to include the whole item set, shorter GIDs indicate larger search spaces for each GID position, and vice versa. We can see that as $l$ varies, the recommendation performance exhibits some fluctuations. For the performance drop from $l=1$ to $l=2$ in Beauty and Sports, the reason is that compared with single token GIDs, GIDs with $l=2$ increase one decoding step but the search space for each step is still large, increasing the search difficulty. When $l=3$, ColaRec achieves a better trade-off between decoding steps and the search space for each step, leading to the best performance in most cases. When $l=4$, a too long GID means more autoregressive decoding steps in model generation, which increases generation difficulty and inference latency. Therefore, we have selected $l=3$ as the default setting in this paper.
To assess how the number of clusters $K$ affects performance, we fix the GID length $l$ as $l=3$ and vary $K$ with values $32$, $64$, $96$, and 128. As shown in Figure 5, a higher $K$ typically results in a slight decrease in overall performance. The decrease in Beauty is more notable. The reason is that a higher $K$ indicates a larger search space in the decoding process, and thus increases the generation difficulty. To this end, choosing a suitable clustering number according to the number of items is important for generative recommendation. The principle is that $K$ needs to be large enough to encode the whole item set, after that $K$ needs to be controlled to limit the search space.
@@ 6. Conclusions and future work
This paper proposes ColaRec, a novel framework to conduct content-based collaborative generation for recommender systems. As an end-to-end generative recommender, ColaRec effectively integrates both item content information and user-item collaborative signals within a unified framework through a generative model. In addition, an auxiliary item indexing task and a contrastive loss are proposed to further align the model’s representations of item content and user-item collaborative signals. We have conducted extensive experiments on four real-world datasets and empirical results demonstrate the effectiveness of ColaRec.
In the future, we intend to investigate more methods to construct generative identifiers (GIDs) and adopt more effective approaches to better align content information and collaborative signals. Negative sampling for generative recommendation is also one of our future works, either using GIDs to sample more informative negative samples or utilizing generative models to generate synthetic negative instances. Besides, how to improve model efficiency for generative recommendation is also a potential research direction.
Acknowledgements
This work was supported by the Tencent WeChat Rhino-Bird Focused Research Program (WXG-FR-2023-07), the Natural Science Foundation of China (62202271, 62372275, 62272274, T2293773,
62102234, 62072279), the National Key R&D Program of China with grant No.2022YFC3303004, the Natural Science Foundation of Shandong Province (ZR2021QF129) and the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001). All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.